Gasoline combusted in a car engine contains 1% (by mole) unburned VOCs. Each gram of gasoline produces 16 grams of exhaust gas. Calculate the concentration of the unburned VOCs in the exhaust gas in ppm. (molecular weight of exhaust gas = 30 g/mol; molecular weight of gasoline = 100 g/mol)
Answer #1

basis : 1 gms of gasoline, moles of gasoline= 1/100= 0.01 g/mole

moles of VOC = 0.01*1/100= 1*10-4 moles

mass of VOC= 1*10-4*30 = 0.003 gms

mass of exhaust gas= 16 gms ( 1 gm of gasolune produces 16 gms of exhasut)

concentration of VOC= 0.003/ (16+0.003)=0.000187= 0.000187*106/106= 187 ppm